### v1

# Think Before You Talk: Enhancing Meaningful Dialogue Generation in Full-Duplex Speech Language Models with Planning-Inspired Text Guidance

Abstract
Full-Duplex Speech Language Models (FD-SLMs) are specialized foundation models designed to enable natural, real-time spoken interactions by modeling complex conversational dynamics such as interruptions, backchannels, and overlapping speech, and End-to-end (e2e) FD-SLMs leverage real-world double-channel conversational data to capture nuanced two-speaker dialogue patterns for human-like interactions. However, they face a critical challenge—their conversational abilities often degrade compared to pure-text conversation due to prolonged speech sequences and limited high-quality spoken dialogue data. While text-guided speech generation could mitigate these issues, it suffers from timing and length issues when integrating textual guidance into double-channel audio streams, disrupting the precise time alignment essential for natural interactions. To address these challenges, we propose TurnGuide, a novel planning-inspired approach that mimics human conversational planning by dynamically segmenting assistant speech into dialogue turns and generating turn-level text guidance before speech output, which effectively resolves both insertion timing and length challenges. Extensive experiments demonstrate our approach significantly improves e2e FD-SLMs’ conversational abilities, enabling them to generate semantically meaningful and coherent speech while maintaining natural conversational flow. [^1] Demos and code will be available at https://dreamtheater123.github.io/TurnGuide-Demo/ and https://github.com/dreamtheater123/TurnGuide.
1 Introduction
Figure 1: Illustration of the two types of FD-SLMs. Cascaded FD-SLMs utilize a state predictor to guide predefined behaviors, while e2e FD-SLMs capture the complex dynamics of real-world double-channel dialogue data.
Speech Language Models (SLMs) have recently attracted significant attention for their role as foundation models in facilitating speech-based interactions with human users (Cui et al. 2024; Hurst et al. 2024). These models leverage advanced autoregressive modeling techniques to produce coherent speech aligned with contextual cues. To facilitate seamless and natural communication experiences, Full-Duplex Speech Language Models (FD-SLMs), a specialized type of SLM, aspire to emulate human-like conversations (Nguyen et al. 2023; Wang et al. 2025). FD-SLMs model user channel and assistant channel audio data simultaneously to support real-time communication characterized by intricate spoken conversational dynamics, including interruptions (Chen et al. 2025), backchannels (Défossez et al. 2024), overlapping speech (Nguyen et al. 2023), and other features that closely mimic natural human interactions. Existing FD-SLMs can be categorized as cascaded and end-to-end (e2e) systems. Cascaded FD-SLMs utilize a separate prediction module to classify spoken dialogues into predefined states, guiding SLM generation (Chen et al. 2025; Wang et al. 2024). While demonstrating efficacy in handling elementary duplex features like interruptions, they are inherently bounded by their dependence on explicit rule-based frameworks. In contrast, e2e FD-SLMs, which we adopt, leverage real-world conversational data to train models end-to-end, capturing nuanced dialogue patterns for more natural, human-like interactions (Wang et al. 2025; Veluri et al. 2024). Figure 1 illustrates these two FD-SLM types. The right side depicts the complex conversational patterns in natural human dialogues, which are modeled by e2e FD-SLMs.
While e2e FD-SLMs facilitate more natural conversational experiences, a significant challenge arises when training these models on spoken dialogue data. Their conversational abilities tend to degrade in comparison to text-based Large Language Models (TLMs) due to the nature of prolonged speech sequences and the scarcity of high-quality speech training data (Zeng et al. 2024; Xie and Wu 2024). One promising solution to this challenge is to transfer the text abilities to speech abilities by leveraging text-guided speech generation, where the model first generates the assistant’s text response based on the input context, followed by generating the corresponding speech response. However, applying text guidance in e2e FD-SLMs poses significant challenges, as these models require precise time alignment between the user and assistant audio streams to effectively capture complex conversational patterns and provide timely, meaningful responses. Integrating text elements into the audio streams can easily disrupt this alignment, ultimately compromising the responsiveness and coherence of the generated speech.
The challenges of integrating texts into assistant-channel audio data can be summarized into two main aspects. 1) Insertion Timing. The text must be inserted precisely at the moment the assistant begins speaking. If the insertion is too early, the assistant may respond without accessing the full context of the user’s speech, leading to misinterpretations or hallucinations. Conversely, if the insertion is too late, the assistant’s speech may precede the inserted text, rendering the text guidance approach ineffective. 2) Insertion Length. The length of the inserted text is critical. If the text is too lengthy, the assistant may continue speaking excessively and struggle to adapt its responses to subsequent user inputs. On the other hand, if the inserted text is too brief, the semantic content may become overly fragmented, diminishing the effectiveness of the text guidance approach (Yao et al. 2025).
To address these challenges, we propose a novel approach, TurnGuide, to seamlessly integrate text data into real-world double-channel spoken conversational data for the training of e2e FD-SLMs. Our approach draws inspiration from the human planning process during conversations, where individuals think ahead, assess real-time speech by others, and plan what to say, when to say it, and for how long. Specifically, we dynamically segment the assistant’s speech into turns, which represent the reasonable lengths of spoken content that convey the assistant’s intended expression, reflecting the structured planning process behind conversational flow. The assistant is trained to plan and output text guidance for each turn before outputting the corresponding speech. By doing so, our method effectively addresses both insertion timing and insertion length challenges. Extensive experiments demonstrate that our approach significantly improves e2e FD-SLMs’ conversational ability, enabling them to generate semantically meaningful and coherent speech while maintaining the natural conversational flow of the generated full duplex dialogue. To summarize our contributions:
  1. 1.
    We propose a novel planning-inspired approach, TurnGuide, which plans both the textual content and the turn-taking timing before speaking, enabling the effective text-guided speech generation for e2e FD-SLMs.
  2. 2.
    Our approach includes a dynamic segmentation and text-guided dialogue modeling framework, addressing the core challenges of insertion timing and length in double-channel dialogues.
  3. 3.
    We demonstrate through extensive experiments that our proposed approach greatly enhances the capability of FD-SLMs to produce semantically meaningful and coherent speech, while preserving the natural conversational flow in full-duplex spoken dialogues.
2 Related Works
Figure 2: Illustration of the multi-modal turn segmentation and alignment framework, which dynamically segments the dialogue speech into turns and aligns the text with speech turns. The text turns serve as the contents derived by the planning process.
2.1 Full Duplex Speech Language Models
Full Duplex Speech Language Models (FD-SLMs) process double-channel conversational data to interact with users seamlessly. FD-SLMs can be categorized into cascaded and end-to-end (e2e) approaches.
Cascaded FD-SLMs
Cascaded FD-SLMs utilize a separate prediction module to classify full-duplex dialogue into a finite number of predefined states, which are typically designed to facilitate transitions between listening and speaking to guide SLM behavior. Training cascaded FD-SLMs consists of two key components: a state predictor, which predicts the SLM’s next state, and a backbone SLM, typically trained on synthetic conversational data. Therefore, cascaded FD-SLMs are constrained by their predefined behaviors, preventing them from authentically engaging in human-like spoken conversations.
The main focus of cascaded FD-SLMs is on defining states. For instance, LSLM (Ma et al. 2025) uses states like ”interrupted” and ”keep speaking”, where interruptions are triggered by specific keywords. Mini-Omni 2 (Xie and Wu 2024) specifies ”start speaking” and ”stop speaking,” halting generation upon hearing ”Stop Omni!”. Freeze-Omni (Wang et al. 2024) bases states on whether the user has finished speaking, prompting the model to wait or interrupt. Minmo (Chen et al. 2025) simplifies its states to whether the model should speak up or keep listening. All these methods rely on synthetic conversational data to manage state transitions, limiting their ability to foster natural, human-like spoken interactions.
End-to-end FD-SLMs
E2e FD-SLMs, on the other hand, employ a transformer-based foundation model architecture (Vaswani et al. 2017; Achiam et al. 2023) to learn from real-world double-channel conversational audio data. This e2e approach enables SLMs to autonomously capture intricate conversational dynamics from authentic human interactions, enabling more natural conversations.
The main focus of e2e FD-SLMs is on effectively modeling double-channel audio data and capturing the interactions between these channels. dGSLM (Nguyen et al. 2023) introduces a Dialogue Language Model that employs distinct transformers to process each audio channel separately, using cross-attention mechanisms to model their interactions. SyncLLM (Veluri et al. 2024) takes a different approach by jointly modeling the double-channel audio within a single LLM, interleaving the two audio sequences into a unified sequence. Moshi (Défossez et al. 2024) adopts an RQ-Transformer architecture (Lee et al. 2022), designed to predict the assistant channel’s output based on the user channel’s input at each time step. NTPP (Wang et al. 2025) proposes a Next-Token-Pair Prediction training objective, enabling simultaneous prediction of audio tokens for both channels in each step.
2.2 Text-Guided Speech Generation in SLMs
SLMs enable end-to-end spoken interactions by generating output speech based on user speech input. However, studies have shown that SLMs exhibit significantly lower performance compared to TLMs counterparts (Cui et al. 2025; Chen et al. 2025; Long et al. 2025). To mitigate this, researchers focus on integrating text to guide the speech generation process, aiming to enhance SLMs by transferring capabilities from TLMs.
Transferring abilities from TLMs to SLMs involves two key processes: text-speech alignment and end-to-end speech interaction. Text-speech alignment ensures the connection between text and speech representations within the SLM backbone, enabling knowledge transfer from text modality to speech modality. This alignment is commonly achieved through Automatic Speech Recognition (ASR) and Text-To-Speech (TTS) tasks during SLM pre-training (Zeng et al. 2024; Wang et al. 2024; Chen et al. 2025). Some approaches also focus on partially replacing speech data with its corresponding text, creating interleaved text-speech sequences for pre-training purposes (Nguyen et al. 2025; Zeng et al. 2024). Following this, SLMs are trained using speech interaction datasets to achieve seamless end-to-end capabilities for speech-based interactions, effectively enhancing their performance and functionality.
The exploration of the text-guided speech generation for e2e FD-SLMs remains in its infancy. Currently, the only notable study in this area is Moshi (Défossez et al. 2024), which proposes adding text tokens corresponding to their respective speech sounds as they occur. However, this method fragments text semantics excessively and does not significantly improve SLM performance. In contrast, our approach adopts a more holistic strategy, planning the overall context before the text insertion process. By systematically inserting text by turns, TurnGuide achieves substantial enhancements in SLM performance.
3 Methodology
Figure 3: Illustration of our text-guided full duplex dialogue modeling framework with two interleaving strategies.
In this section, we present the comprehensive methodology of TurnGuide, designed to enhance the e2e FD-SLM by utilizing text guidance. This approach is divided into two key frameworks: one for planning and segmenting spoken dialogue into turns and performing speech-text alignment, and the other for modeling text-guided full-duplex dialogue.
3.1 Multi-Modal Turn Segmentation and Alignment Framework
To enable text guidance in full-duplex speech interactions, it is essential to derive training data that effectively combines text and speech. As a first step, it is crucial to determine how text can be incorporated without disrupting the natural conversational flow. This requires the text to be introduced at an appropriate moment and kept within a reasonable length. We propose tackling these challenges by segmenting the speech into turns and aligning each part of the text with its corresponding speech turn.
We leverage both speech and text as multi-modal cues to perform turn segmentation, as shown in Figure 2. We first apply Voice Activity Detection (VAD) to identify speech segments within the audio stream. Let $A$ represent the input audio sequence, VAD produces a list of speech segments:
$V=\text{VAD}(A)=\{(\tau_{1}^{start},\tau_{1}^{end}),(\tau_{2}^{start},\tau_{2}^{end}),$...\}, (1)
where each tuple $(\tau_{i}^{start},\tau_{i}^{end})$ represents the start and end timestamps of the $i$-th detected speech segment. We use the pyannote library (Bredin et al. 2020) as the VAD module. Considering the VAD result is too fragmented, we merge nearby speech segments into Inter-Pausal Units (IPUs) (Nguyen et al. 2023), which are defined by consecutive segments having pauses within a threshold $\tau_{IPU}=0.5$ seconds, i.e.,
$IPU_{j}=\bigcup_{i\in G_{j}}(\tau_{i}^{start},\tau_{i}^{end}),$ (2)
where $G_{j}=\{i:\tau_{i+1}^{start}-\tau_{i}^{end}<\tau_{IPU}\}$ defines the group of consecutive segments to be merged into IPU $j$.
To further refine the turn segmentation, we utilize text transcripts with word-level timestamps aligned to the speech audio. While some speech dialogue datasets include transcriptions, we use an Automatic Speech Recognition (ASR) module with timestamp functionality to perform the task. This approach ensures adaptability to unlabeled dialogue datasets. The ASR process is formulated as
$W=\text{ASR}(A)=\{w_{1},w_{2},$...\}, (3)
where $w_{i}$ represents the $i$-th word in the transcription. The word-level timestamp extraction process is formulated as
$\Theta=\text{TE}(A)=\{(\delta_{1}^{start},\delta_{1}^{end}),(\delta_{2}^{start},\delta_{2}^{end}),$...\}, (4)
where $(\delta_{i}^{start},\delta_{i}^{end})$ represents the start and end timestamps of word $w_{i}$. In practice, we employ the Whisper medium model (Radford et al. 2023) for both ASR and timestamp extraction tasks. To enhance the stability of the timestamp extraction process, we utilize the whisper-timestamped package (Louradour 2023).
Finally, we utilize timestamp information to enhance turn segmentation further. First, we group each word into its associated IPU using $\tau^{start}$. We incorporate a tolerance parameter $\tau_{tol}=0.6$ seconds to adjust the word segmentation boundary slightly. This adjustment accounts for instances where, in practice, the first word of an IPU might incorrectly align with the preceding IPU due to inconsistencies in timestamping. The process is defined as
$W_{j}=\{w_{i}:\tau_{IPU_{j}}^{start}-\tau_{tol}\leq\delta_{i}^{start}\leq\tau_{IPU_{j}}^{end}-\tau_{tol}\},$ (5)
where $W_{j}$ represents words grouped into turn $j$, and $\tau_{IPU_{j}}^{start}$ and $\tau_{IPU_{j}}^{end}$ represent the start and end times of IPU $j$. We denote the speech segment for IPU $j$ as $S_{j}$. Then, we split both the grouped text segments and their corresponding speech segments at natural punctuation boundaries to create coherent dialogue turns. Each word group $W_{j}$ is segmented at punctuation marks $P=\{$,”, “.”, “??”, “!!}\}, i.e., $W_{j}^{final}=\text{split}(W_{j},P)$, and we extract the corresponding speech segment $S_{j}^{final}$ according to the word-level timestamps. The final output consists of aligned text-speech pairs $\{(W_{j}^{final},S_{j}^{final})\}$ where each text segment is temporally aligned with its corresponding speech segment, serving as training data for text-guided full-duplex dialogue modeling.
3.2 Text-Guided Full Duplex Dialogue Modeling
After deriving the aligned dialogue turns, the next step is to model double-channel audio data with text guidance. We use GLM-4-Voice (Zeng et al. 2024), an off-the-shelf SLM, as the backbone for training on full-duplex dialogue datasets. GLM-4-Voice includes a speech tokenizer to convert audio into discrete tokens, a language model to process these tokens, and a vocoder to reconstruct speech. Its strong speech-text alignment makes it ideal for the backbone of text-guided speech generation. As illustrated in Figure 3, our text-guided full duplex dialogue modeling framework centers around two interleaving strategies:
Channel-wise Interleaving. Full-duplex dialogue modeling necessitates the simultaneous processing of two-channel audio to capture cross-channel interactions. A common technique to address this is by interleaving the two audio channels into a single sequence, either with token-level interleaving (Wang et al. 2025) or chunk-level interleaving (Veluri et al. 2024; Zhang et al. 2024). We choose chunk-level over token-level interleaving because it demonstrates better performance in our experiments. We first consider channel-wise interleaving for speech tokens. To implement this technique, let $s_{u}$ and $s_{a}$ represent speech tokens from the user and assistant channels. Each channel’s speech tokens are divided into speech chunks with a chunk size of $\tau_{SC}=5$. Given GLM-4-Voice’s frame rate of $fr=12.5\text{Hz}$, each speech chunk corresponds to a 400-ms audio segment. The chunk-interleaving process alternates segments between the two channels, forming a sequence $SC=[SC_{u}^{1},SC_{a}^{1},SC_{u}^{2},SC_{a}^{2},\dots]$. Each speech chunk contains $\tau_{SC}$ speech tokens, with a unique speaker ID token inserted at the beginning to distinguish between the channels. Specifically, $SC_{u}^{i}=[UID,s_{u}^{i\times\tau_{SC}+1},s_{u}^{i\times\tau_{SC}+2},\dots,s_{u}^{i\times\tau_{SC}+\tau_{SC}}]$ and $SC_{a}^{i}=[AID,s_{a}^{i\times\tau_{SC}+1},s_{a}^{i\times\tau_{SC}+2},\dots,s_{a}^{i\times\tau_{SC}+\tau_{SC}}]$ represent the $i$-th user and assistant channel speech chunks, where $UID$ and $AID$ denote the user and assistant ID tokens.
Text-speech Interleaving. We apply the text-guided speech generation strategy on the assistant channel. Specifically, we interleave the text tokens with speech tokens for each turn, starting from the first speech chunk of the turn. Given the text in a turn $W_{j}^{final}$, we first append an “end of text” ($\langle EOT\rangle$) token at the end of it (i.e., $W_{j}^{final}=[t_{j}^{1},t_{j}^{2},...,\langle EOT\rangle]$) and then split it into text chunks $TC_{j}=[TC_{j}^{1},TC_{j}^{2},$...] with chunk size $\tau_{TC}=5$, where $TC_{j}^{i}=[t_{j}^{i\times\tau_{TC}+1},t_{j}^{i\times\tau_{TC}+2},...,t_{j}^{i\times\tau_{TC}+\tau_{TC}}]$ represents the $i$-th text chunk for turn $j$. Second, we calculate the first speech chunk for text-speech interleaving for $W_{j}^{final}$ as
$k_{j}^{start}=\lfloor\frac{\tau_{j}^{start}\times fr}{\tau_{SC}}\rfloor.$ (6)
Then, the interleaving is performed from $k_{j}^{start}$ until the last text chunk as $TC_{j}^{0},SC_{a}^{k_{j}^{start}},SC_{u}^{1},SC_{a}^{k_{j}^{start}+1},$...].
The final sequence is formed by chunks with the two interleaving strategies, $FC=[C_{u}^{1},C_{a}^{1},C_{u}^{2},C_{a}^{2},$...], where
$C_{u}^{i}=SC_{u}^{i}\text{ and }C_{a}^{i}=\begin{cases}[TC_{j}^{k},SC_{a}^{i}]&\text{if text chunk presents}\\ SC_{a}^{i}&\text{otherwise}.\end{cases}$
The decision to use a 5:5 ratio for text chunks and speech chunks ensures that text generation precedes speech generation, providing effective text guidance. In summary, our full duplex dialogue modeling framework leverages two interleaving strategies to enable both cross-channel interaction and text-guided speech modeling.
SLMs Fisher-finetuned Open-source
dGSLM (Nguyen et al. 2023)
SyncLLM (Veluri et al. 2024)
Moshi (Défossez et al. 2024)
OmniFlatten (Zhang et al. 2024)
NTPP (Wang et al. 2025)
Wu et al. (Wu et al. 2025)
Table 1: Summary of the open-source status of previous e2e FD-SLMs papers.
SLM Semantics Turn-taking (Pearson Correlation)
GPT-score Occ CumDur AvgDur Overall
dGSLM 1.32 0.31 0.24 0.13 0.23
STI 2.55 0.32 0.29 0.14 0.25
SCI 2.71 0.31 0.35 0.17 0.28
Table 2: Comparison of both the semantic (GPT-score, measured on Fisher dataset) and turn-taking abilities (measured on Candor dataset) between our trained baselines and dGSLM. Occ, CumDur, and AvgDur represent occurrence, cumulative duration, and averaged duration, respectively.
SLM Unconditional Conditional - All Conditional - Assistant Overall
0.8 0.9 1.0 1.3 AVG 0.8 0.9 1.0 1.3 AVG 0.8 0.9 1.0 1.3 AVG
SCI 1.81 2.17 2.48 2.71 2.29 1.70 1.87 1.98 2.16 1.93 1.35 1.52 1.67 2.00 1.63 1.95
Moshi TS 1.72 2.06 2.41 2.74 2.23 1.46 1.63 1.82 2.13 1.76 1.23 1.38 1.57 1.96 1.54 1.84
TurnGuide 2.68 2.89 2.93 2.84 2.84 2.08 2.15 2.25 2.26 2.19 2.02 2.12 2.17 2.24 2.14 2.39
TurnGuide L2:1 2.87 3.06 3.08 2.89 2.98 2.20 2.32 2.38 2.35 2.31 2.18 2.33 2.41 2.35 2.32 2.54
TurnGuide L3:1 2.94 3.05 3.08 2.86 2.98 2.25 2.31 2.33 2.32 2.30 2.27 2.38 2.38 2.37 2.35 2.54
Performance Upper Bounds
TurnGuide GT Text 2.61 2.77 2.91 2.82 2.78 2.22 2.32 2.36 2.30 2.30 2.10 2.29 2.30 2.30 2.25 2.44
TextOnly 3.30 3.30 3.28 2.82 3.18 - - - - - - - - - - -
Table 3: Semantic evaluation results (0-5) of different FD-SLMs under different temperature settings by GPT-4o.
Figure 4: Perplexity evaluation of FD-SLMs across temperatures using Fisher-text-only, GLM-4-Voice, and Llama-3.1-8B-Instruct as evaluation models.
4 Experiments
In this section, we assess the effectiveness of TurnGuide, primarily focusing on the semantic meaningfulness of the generated dialogues.
4.1 Experimental Setups
Datasets
We mainly utilize the Fisher dataset (Cieri, Miller, and Walker 2004), a double-channel telephone conversation corpus containing 2,000 hours of audio data, for training and evaluating our FD-SLMs. Each recording in the Fisher dataset spans approximately 10 minutes, which we segment into 120-second clips for training and evaluation. We divide the Fisher dataset into training, validation, and test sets, with detailed information provided in Appendix E. Additionally, we incorporate the Candor dataset (Reece et al. 2023), another double-channel conversation corpus, for evaluation purposes.
Baselines and Our Approach
We consider the following models or methods as our baselines:
  • dGSLM (Nguyen et al. 2023): a dual-Transformer architecture, each modeling an individual audio channel.
  • Speech Token Interleaving (STI): We train the GLM-4-Voice backbone on the Fisher dataset by interleaving the speech tokens from the two channels, one token at a time.
  • Speech Chunk Interleaving (SCI): Similar to Speech Token Interleaving, but it interleaves speech tokens chunk by chunk with a chunk size of 5.
  • Text-speech Token Alignment (Moshi Training Strategy, Moshi TS (Défossez et al. 2024)): The only existing method for text-guided speech generation in e2e FD-SLMs. It involves inserting each text token exactly at the time its corresponding sound is produced. However, Moshi is trained using a wide range of datasets, with Fisher being only one of them, making it not directly comparable to our Fisher-finetuned baselines. To address this, we train the GLM-4-Voice backbone on the Fisher dataset with Moshi’s training methodology to compare it with other baselines. Our detailed implementation of Moshi TS is illustrated in Appendix C.
  • Our Approach (TurnGuide): We train the GLM-4-Voice backbone on the Fisher dataset by utilizing the text-guided full duplex dialogue modeling framework introduced in Section 3.2. The text-speech dataset is pre-processed using the framework illustrated in Section 3.1.
Remark 1
While there are other e2e FD-SLMs apart from dGSLM, most have not been (fully) open-sourced yet, as summarized in Table 1. Therefore, to evaluate our proposed text guidance approach, we train baseline models and show that they perform reasonably well. We then compare TurnGuide against our trained baselines.
Remark 2
Most e2e FD-SLMs demonstrate effectiveness by showing superiority over dGSLM (Veluri et al. 2024; Wang et al. 2025; Défossez et al. 2024). Therefore, we use the same way to demonstrate that our trained baseline performs reasonably well.
Model Training and Inference
The models are trained on the Fisher dataset for 2 epochs on 8 A800 GPUs with a global batch size of 256, a learning rate of 4e-6, a cosine learning rate scheduler, no weight decay, and 20 warmup steps. During inference, we sample 1,000 audio clips from the Fisher and Candor test sets. The first 30 seconds serve as prompts, while the model generates 90-second continuations using two methods: unconditional and conditional generation. In unconditional generation, we train the model with text guidance in both audio channels, and the model generates both channels for the 90-second continuation during inference, whereas in conditional generation, we train the model with only assistant-channel text guidance, and the model generates the assistant-channel audio based on the user-channel audio as context.
SLM Occurrence Cumulative Duration Averaged Duration Overall
0.8 0.9 1.0 1.3 AVG 0.8 0.9 1.0 1.3 AVG 0.8 0.9 1.0 1.3 AVG
Unconditional
SCI 0.57 0.57 0.61 0.62 0.59 0.48 0.48 0.52 0.58 0.51 0.26 0.25 0.33 0.36 0.30 0.47
Moshi TS 0.52 0.53 0.56 0.61 0.56 0.38 0.43 0.46 0.54 0.45 0.20 0.24 0.28 0.34 0.26 0.42
TurnGuide 0.57 0.57 0.63 0.62 0.60 0.46 0.52 0.53 0.56 0.52 0.27 0.25 0.34 0.36 0.30 0.47
TurnGuide L2:1 0.57 0.60 0.62 0.62 0.60 0.50 0.51 0.55 0.56 0.53 0.29 0.26 0.36 0.37 0.32 0.48
TurnGuide L3:1 0.58 0.59 0.63 0.58 0.60 0.47 0.50 0.54 0.55 0.51 0.31 0.25 0.33 0.37 0.32 0.47
Conditional
SCI 0.47 0.49 0.49 0.50 0.49 0.28 0.40 0.46 0.53 0.42 0.17 0.25 0.29 0.34 0.26 0.39
Moshi TS 0.48 0.48 0.48 0.45 0.47 0.39 0.39 0.43 0.50 0.43 0.24 0.25 0.29 0.29 0.27 0.39
TurnGuide 0.46 0.50 0.51 0.46 0.48 0.38 0.46 0.48 0.50 0.45 0.24 0.26 0.34 0.32 0.29 0.41
TurnGuide L2:1 0.50 0.48 0.51 0.43 0.48 0.39 0.49 0.54 0.49 0.48 0.27 0.33 0.33 0.33 0.31 0.42
TurnGuide L3:1 0.46 0.48 0.49 0.47 0.48 0.39 0.46 0.51 0.50 0.46 0.20 0.30 0.28 0.32 0.27 0.41
Table 4: Pearson correlation results of diverse turn-taking events with conditional generation of various FD-SLMs evaluated under different temperature configurations.
4.2 Results
We aim to answer two critical Research Questions (RQs) through comprehensive experiments. RQ1: Does our method significantly improve the semantic meaningfulness and coherence of the generated full-duplex dialogue? RQ2: Can our method preserve the natural conversational flow after integrating additional texts into the sequence?
RQ1
To evaluate the semantics of the generated dialogue, the most intuitive method is to conduct human evaluations. However, considering the time, cost, and potential inconsistencies associated with human evaluations, we propose utilizing the GPT-score. Specifically, given a generated double-channel dialogue audio, we first transcribe the audio from each channel into text with sentence-level timestamps. For this process, we employ the stable-ts package (Jian 2023) with the Whisper-large-v3 model. Subsequently, the transcribed sentences from both channels are grouped based on their start times, effectively representing the dialogue between the two speakers in text form along the timeline. Finally, the consolidated text is analyzed using GPT-4o (Hurst et al. 2024), which assesses the meaningfulness and coherence of the dialogue on a scale of 0 to 5 with detailed rubrics. The complete GPT-4o prompts for unconditional and conditional generation settings are shown in Appendix G.
Baseline Comparison. We compare our trained baselines, STI and SCI, with dGSLM in the unconditional generation setting on the Fisher test set (the same setting used in the dGSLM paper). As shown in Table 2, STI and SCI perform significantly better than dGSLM, showing that our baselines are trained reasonably well.
Effectiveness of Our Approach. We then demonstrate the effectiveness of TurnGuide through evaluations in both unconditional and conditional settings. In the ”Conditional - All” scenario, GPT-4o uses ordered transcripts from both channels to assess responses based on user speech. However, we observed that GPT-4o overly prioritizes user-assistant interactions, neglecting the coherence of assistant-generated speech within its channel. To address this, the ”Conditional - Assistant” setting provides only the assistant-channel transcript, enabling a focused analysis of semantic quality within the assistant channel.
We compare TurnGuide with SCI and Moshi TS under different temperatures (as shown in Table 3) and make the following observations. 1) Significant performance improvements over baselines. TurnGuide performs significantly better than SCI and Moshi TS (over 30% performance gain), showing that TurnGuide successfully leverages the text information to boost the meaningfulness and coherence of the generated dialogue. 2) Speech-only methods and text-speech methods excel at different temperatures. SCI consistently performs the best at the largest temperature (1.3), whereas TurnGuide tend to perform the best at temperature 1.0. Additionally, TurnGuide shows more performance gain under lower temperatures, such as 0.8 and 0.9. 3) Focusing more on text generation makes the model perform better. In “TurnGuide L2:1” and “TurnGuide L2:1” settings, we set the training loss weight of text and speech tokens to be 2:1 and 3:1, and the results show that they consistently outperform the standard 1:1 setting (“TurnGuide”). 4) TurnGuide achieves performance comparable to its upper bounds. In Table 3, “TurnGuide GT Text” refers to the GLM-4-Voice model trained with our proposed framework but switches to the ground truth transcriptions from the Fisher dataset. TurnGuide performs similarly with TurnGuide GT Text in the 1:1 loss setting and outperforms in the 2:1 and 3:1 loss settings. Additionally, “TextOnly” represents a model trained exclusively on transcriptions without speech tokens, where TurnGuide delivers close performance (2.98–3.18), highlighting the effectiveness of our approach.
Further validation with perplexity. To facilitate the semantic evaluation using GPT-score, perplexity is employed to further demonstrate the superior semantic abilities of TurnGuide. Specifically, we calculate perplexity using the same ordered transcript used in the GPT-score evaluation. Perplexity is measured with three distinct language models: Fisher-text-only, GLM-4-Voice, and Llama-3.1-8b-Instruct (Dubey et al. 2024). The Fisher-text-only model, trained using Fisher-provided text transcript data on the GLM-4-Voice base model, serves as the ground truth for text-based dialogue distribution. GLM-4-Voice acts as the SLM backbone, while Llama-3.1-8b-Instruct serves as the text-only model backbone. We analyze the perplexity of generated dialogues across different temperature settings in both conditional and unconditional scenarios. As illustrated in Figure 4, TurnGuide consistently achieves lower perplexity compared to baseline approaches, demonstrating superior semantic modeling abilities.
RQ2
To evaluate whether the full-duplex dialogue generated by TurnGuide preserves a natural conversational flow, we measure turn-taking events (Nguyen et al. 2023), categorized into IPU, pause, gap, and overlap, as defined in (Nguyen et al. 2023), with detailed explanations in Appendix H. We calculate the Pearson correlation of the occurrence, cumulative duration, and average duration of the turn-taking events between the 30-second prompts and the corresponding 90-second continuations, as described in (Veluri et al. 2024; Nguyen et al. 2023). This measures whether the model can generate dialogues with similar conversational features to the prompt. The metric is averaged across the four turn-taking events.
Baseline Comparison. We compare our trained baselines, STI and SCI, with dGSLM in the unconditional generation setting (as employed in the dGSLM paper) using the Candor test set. The Candor test set allows for a fairer comparison, addressing the issue where previous papers did not disclose their test set split from the Fisher dataset, leading to significant test set leakage during evaluation. As presented in Table 2, both STI and SCI outperform dGSLM, showing that our baselines are trained reasonably well.
Preservation of the natural conversational flow. We use the same turn-taking evaluation method and compare TurnGuide with SCI and Moshi TS under different temperatures. We evaluate in both unconditional and conditional settings. As shown in Table 4, TurnGuide shows nearly identical performance in Pearson correlation for occurrence, cumulative duration, and averaged duration, showing TurnGuide successfully preserves the natural conversation flow.
5 Conclusion
In this paper, we present a planning-inspired text-guided approach, TurnGuide, to improve the dialogue generation abilities of end-to-end Full-Duplex Speech Language Models. By segmenting assistant speech into turns and training the model with turn-level text-guided speech generation with channel-wise and text-speech interleaving strategies, our approach effectively addresses key challenges of text insertion timing and length, ensuring coherent and natural dialogue generation. Experiments show significant improvements of our method in semantic quality while maintaining natural conversational dynamics, validating its effectiveness for spoken dialogue systems and paving the way for more meaningful and human-like speech interactions.
Appendix A Limitations and Future Works
While the paper introduces TurnGuide, a novel planning-inspired text-guided speech generation approach for enhancing e2e FD-SLMs, it still presents several limitations. First, the effectiveness of TurnGuide is primarily validated on the Fisher and Candor datasets, which, although widely used, may not fully capture the diversity and complexity of real-world conversational scenarios since both Fisher and Candor are telephone conversation datasets. Second, the evaluation relies heavily on GPT-4o-based automated metrics (GPT-score) for semantic assessment, which, despite being aligned with human judgments in most cases, may not perfectly reflect the nuanced human perceptions of dialogue quality.
We then suggest several promising directions to expand upon TurnGuide. One key avenue is to explore the generalization of the proposed framework to more diverse and challenging conversational datasets, including those featuring multiple speakers, varied languages, or more spontaneous dialogue styles. Enhancing robustness to noisy or low-resource settings—such as improving alignment without high-quality transcriptions or adapting to imperfect voice activity detection—would also broaden real-world applicability. Moreover, integrating additional modalities (e.g., visual cues or gestures) could further enrich dialogue modeling and understanding.
Figure 5: Illustration of the Moshi training strategy.
Appendix B Judgement Alignment between Human and GPT-4o in Semantic Evaluation
Difference 1 2 3 Overall
Accuracy 7/10 9/10 9/10 25/30
Table 5: Alignment accuracy between human judgments and GPT-4o semantic evaluation scores for generated dialogue pairs with varying score differences.
Figure 6: Training and validation loss comparison of different training strategies.
Figure 7: Illustration of the four turn-taking events used in the evaluation.
In this paper, we primarily utilize GPT-4o to evaluate the semantics of the generated dialogues. This raises a crucial question: Can AI evaluation scores accurately capture human preferences? To investigate this, we conduct a human listening test to examine the alignment between GPT-evaluated scores and human judgments. For this experiment, we randomly select 30 pairs of generated dialogues that vary in their GPT scores. Specifically, we group these pairs into three categories: 10 pairs with a GPT-score difference of 1, 10 pairs with a difference of 2, and 10 pairs with a difference of 3. Human judges are then asked to evaluate which dialogue in each pair demonstrates better semantics. To ensure unbiased evaluation, the order of dialogues within each pair is randomized. Additionally, judges are provided with both the audio of the dialogues and their transcripts, generated using Whisper-large-v3 with the stable-ts package. We recruit three human judges, all of whom have native English proficiency, to carry out the evaluations.
The results highlight strong human-AI alignment in this task, as presented in Table 5. Notably, human judgments match GPT-4o in 25 out of the 30 pairs evaluated. Among the mismatches, most occur in cases with minor differences (1-difference), but the judgments still match for 7 out of 10 pairs. This suggests GPT excels at distinguishing dialogues with significant semantic differences, but it may struggle when dialogues exhibit similar semantic levels. These findings align closely with human evaluations, underscoring the reliability of our GPT-score evaluation methodology.
Appendix C Moshi TS Implementation
As detailed in Section 4.1, we do not directly compare our model with Moshi, as Moshi is a general-purpose SLM that utilizes different pre-training speech data and is not fine-tuned on the Fisher dataset during its final training stage. Instead, we opt to train a model using Moshi’s training strategy, employing GLM-4-Voice on the Fisher dataset. This approach ensures a fair comparison, as the base model and training dataset remain identical, with the only difference between our approach and Moshi TS being the training strategy used. Therefore, in this section, we illustrate the details of the training process of Moshi TS.
First, Moshi utilizes eight codebooks, where every time step includes eight speech tokens for each channel. We adapt this technique for the GLM-4-Voice backbone, which relies on a single codebook method, allotting one speech token per channel per time step. Then, the tokens from each channel are interleaved at each time step, making Moshi rely on a token-level interleaving approach. In the original implementation of Moshi, a text-channel token precedes each assistant-channel speech token, which could either be a genuine text token, a text pad token, or an end-of-text pad token. Moshi also employs text-speech interleaving at token-level, where each text token is incorporated at the precise moment it is spoken. Consequently, if a time step features an audible text token, the corresponding text token is inserted. If not, either a text pad token or an end-of-text pad token is inserted, determined by whether the subsequent text-channel token is an actual text token. We follow the original Moshi implementation by extracting the timestamp of each word with Whisper-timestamped package (Louradour 2023) and the medium Whisper model (Radford et al. 2023). The illustration of our implemented Moshi training strategy is shown in Figure 5.
Appendix D Comparison of Different Training Strategies
In this section, we provide a comparison of different training strategies and share corresponding insights. We start by examining the impact of these strategies on the training process by analyzing the training loss curves displayed in Figure 6. First, SCI achieves a significant reduction in training loss compared to STI. This improvement can likely be attributed to the addition of speaker ID tokens for each channel, which provides extra information about the correspondence between speech tokens and their respective speakers, thereby simplifying the model’s learning process. Furthermore, the reduction in training loss may also result from the organization of speech tokens into chunks, which enhances contextual coherence compared to STI. Second, TurnGuide demonstrates even lower training loss than STI. More specifically, TurnGuide outperforms both SCI and STI, with configurations such as TurnGuide L2:1 and TurnGuide L3:1 achieving even greater reductions in training loss. These results highlight the benefits of the text-speech alignment process, which facilitates SLM training and enables superior semantic performance. Finally, the loss reduction observed in Moshi TS does not translate into performance improvements. While Moshi TS achieves one of the lowest training losses, its overall performance trails behind our approach. This discrepancy may stem from the excessive presence of text pad tokens relative to actual text tokens. The repetitive nature of pad tokens simplifies prediction, leading to artificially low training losses without corresponding gains in real-world performance.
We then intuitively explore the reasons behind the lower performance of Moshi TS compared to TurnGuide and identify three key factors contributing to this difference. First, Moshi TS tends to fragment text semantics. Both Moshi TS and our method aim to enhance the semantics of generated speech using textual information. However, Moshi TS splits text tokens so significantly that it hampers the model’s ability to learn coherent text representations. In contrast, TurnGuide aggregates these text tokens, creating a more cohesive grouping that simplifies the learning process for text generation, which leads to notable performance improvements. Second, Moshi TS fails to mitigate the inherently fragmented nature of speech semantics. Here, we propose a new way of understanding the semantic information within a speech audio. Speech audio combines multiple types of information together, including semantic, paralinguistic, and speaker information (Cui et al. 2024), and the semantics within speech can be defined by its transcription, with speech containing semantic information only at the temporal points when each word is sounded. This fragmented nature increases the difficulty for coherent speech generation. Although Moshi TS incorporates text into the speech generation process, it inserts text tokens strictly at the time they are sounded in the speech. This approach only supplements speech with semantic information at points where semantic information already exists, offering limited potential for enhancing semantic representation. By comparison, TurnGuide groups texts within the same turn and processes them collectively. This method introduces denser semantic information, effectively guiding the speech generation process to improve coherence. Third, Moshi TS heavily depends on precise word-level timestamps from the ASR systems. In practice, these timestamps are not always exact, often being slightly ahead or behind the actual sounding moment, which can disrupt text-speech alignment and degrade performance. TurnGuide, however, leverages only the start time of each turn, significantly reducing reliance on word-level timestamp accuracy. This simpler alignment strategy not only lowers the risk of errors but also contributes to stronger overall performance.
Appendix E Fisher and Candor Dataset Split
Both the Fisher and Candor datasets are frequently referenced and employed in e2e FD-SLM studies. However, to date, there is no official training, validation, and test set split for these datasets, nor do existing papers provide details of their dataset splits. This poses a significant challenge for future research, as it becomes difficult to replicate the reported model performance in original studies. Specifically, when researchers use their own dataset splits to train models and subsequently compare their model performance with publicly available checkpoints from prior work, there is a risk of test set leakage. This occurs when an existing model has been trained on test samples included within a new split defined by researchers. To tackle this issue, we propose releasing our train/validation/test splits for the Fisher and Candor datasets and strongly encourage future studies to adopt these splits for a more equitable and reliable basis of comparison. Specifically, Fisher separates all the audio into 117 parts (000-116). We split the train/validation/test set using part IDs 000-110/111-113/114-116, and we randomly select roughly 1,000 samples in our Fisher test set to perform all the evaluations reported in this paper. For the Candor dataset, we randomly select roughly 1,000 samples from the part 001 audio data to perform the evaluations. We will report the IDs of the 1,000 samples for both datasets after the acceptance of the paper.
Appendix F Detailed Computing Infrastructure Specification
We perform all the experiments on A800 GPUs, which have 80GB of GPU memory. The CPU model of the machine is Intel(R) Xeon(R) Platinum 8378A CPU @ 3.00GHz. The operating system is Ubuntu 20.04.4 LTS. Our training environment uses Pytorch 2.0.0 with the transformers library of version 4.49.0. Our inference environment uses Pytorch 1.5.0 with the transformers library of version 4.44.2. We run each experiment once.
Appendix G Semantic Evaluation Prompts
We give prompts for the three settings of the GPT-score semantic evaluation used in Table 3: We use the prompt in Figure 8 for the unconditional setting, and use the prompt in Figure 9 for the two conditional settings.
Appendix H Turn-taking Events Explanations
This section details the four turn-taking events utilized in our evaluation: Inter-Pausal Units (IPUs), Pause, Gap, and Overlap. Figure 7 provides examples of these turn-taking events. First, the speech segments within each speaker channel are identified using a Voice Activity Detection (VAD) module. An IPU is defined as continuous speech within a single speaker channel. Multiple speech segments are grouped into a single IPU if the pause between them is shorter than $\tau_{\text{IPU}}$. As previously mentioned, $\tau_{\text{ IPU }}$ is initially set to 0.5 seconds in the multi-modal turn segmentation and alignment framework (Section 3.1). However, for the evaluations, it is adjusted to 0.2 seconds in accordance with prior studies (Nguyen et al. 2023; Veluri et al. 2024; Wang et al. 2025). A Pause refers to the absence of speech within the same speaker channel, while a Gap indicates the discontinuation of speech across different channels. Overlap refers to the time period during which both speakers are speaking simultaneously.
Please evaluate the following two-speaker dialogue transcript for how meaningful the speech is (based on its content), focusing primarily on the part from 30 seconds to 120 seconds of the transcript. Use the following scale:
0: Completely meaningless; no coherent sentences, random words, or unintelligible.
0.5: Almost no meaning; isolated words or phrases, but no understandable ideas.
1: Extremely low meaning; rare, vague fragments of ideas, but mostly incoherent or off-topic.
1.5: Very little meaning; a few short, unclear ideas, but mostly disjointed or confusing.
2: Low meaning; some recognizable ideas or topics, but mostly unclear, incomplete, or off-topic.
2.5: Somewhat low meaning; a few coherent points, but overall lacks clarity or logical flow.
3: Moderate meaning; general topic is understandable, but there are gaps, unclear parts, or weak connections.
3.5: Fairly meaningful; mostly coherent and relevant, but with some confusion, repetition, or lack of detail.
4: Meaningful; clear and logical, with relevant and connected ideas, though may lack depth or detail.
4.5: Very meaningful; almost fully coherent, with well-developed, relevant, and connected ideas.
5: Extremely meaningful; highly coherent, clear, and detailed, with all ideas well connected and relevant.
Only output the final score (0, 0.5, 1, 1.5, ..., 5) **ONLY** according to the above rubric. Do not output anything else.
Figure 8: The prompt for GPT-score semantic evaluation under the unconditional setting.
Please evaluate the following two-speaker dialogue transcript for how meaningful the speech is (based on its content), only focusing on the model channels output and from 30 seconds to 120 seconds of the transcript. Use the following scale:
0: Completely meaningless; no coherent sentences, random words, or unintelligible.
0.5: Almost no meaning; isolated words or phrases, but no understandable ideas.
1: Extremely low meaning; rare, vague fragments of ideas, but mostly incoherent or off-topic.
1.5: Very little meaning; a few short, unclear ideas, but mostly disjointed or confusing.
2: Low meaning; some recognizable ideas or topics, but mostly unclear, incomplete, or off-topic.
2.5: Somewhat low meaning; a few coherent points, but overall lacks clarity or logical flow.
3: Moderate meaning; general topic is understandable, but there are gaps, unclear parts, or weak connections.
3.5: Fairly meaningful; mostly coherent and relevant, but with some confusion, repetition, or lack of detail.
4: Meaningful; clear and logical, with relevant and connected ideas, though may lack depth or detail.
4.5: Very meaningful; almost fully coherent, with well-developed, relevant, and connected ideas.
5: Extremely meaningful; highly coherent, clear, and detailed, with all ideas well connected and relevant.
Only output the final score (0, 0.5, 1, 1.5, ..., 5) **ONLY** according to the above rubric. Do not output anything else.
Figure 9: The prompt for GPT-score semantic evaluation under the two conditional settings.