### v2

# TurnGuide: Enhancing Meaningful Full Duplex Spoken Interactions via Dynamic Turn-Level Text-Speech Interleaving

Abstract
Full-Duplex Speech Language Models (FD-SLMs) are specialized foundation models designed to enable natural, real-time spoken interactions by modeling complex conversational turn-taking such as interruptions, backchannels, and overlapping speech. End-to-end (e2e) FD-SLMs leverage real-world double-channel conversational data to capture nuanced two-speaker dialogue patterns for human-like interactions, but their conversational abilities often degrade compared to pure-text conversation due to prolonged speech sequences and limited high-quality spoken dialogue data. Although interleaved text-speech generation could mitigate this degradation, integrating discrete text tokens into continuous double-channel audio streams could disrupt the precise time alignment required for fluid interaction. To address this, we propose TurnGuide, a novel text-speech interleaved generation approach for e2e FD-SLMs that dynamically segments assistant speech into dialogue turns and interleaves turn-level text and speech generation. This approach allows FD-SLMs to integrate the semantic intelligence of LLMs without compromising the natural acoustic flow. Extensive experiments show that TurnGuide not only significantly improves e2e FD-SLMs to produce semantically meaningful, coherent speech but also achieves state-of-the-art performance on various turn-taking events. [^1] Demos and code will be available at https://dreamtheater123.github.io/TurnGuide-Demo/ and https://github.com/dreamtheater123/TurnGuide.
TurnGuide: Enhancing Meaningful Full Duplex Spoken Interactions via Dynamic Turn-Level Text-Speech Interleaving
Wenqian Cui1thanks: Project lead. Email: wenqian.cui@link.cuhk.edu.hk, Lei Zhu2, Xiaohui Li2thanks: Corresponding author., Zhihan Guo1, Haoli Bai2, Lu Hou2, Irwin King122footnotemark: 2 1The Chinese University of Hong Kong, 2Huawei Technologies
1 Introduction
Figure 1: Illustration of the two types of FD-SLMs. Cascaded FD-SLMs utilize a state predictor to guide predefined behaviors, while e2e FD-SLMs capture the complex dynamics of real-world double-channel dialogue data.
Speech Language Models (SLMs) have attracted attention as foundation models for speech-based interactions with human users Cui et al. (2024); Hurst et al. (2024). Full-Duplex Speech Language Models (FD-SLMs), a specialized type of SLM, model user- and assistant-channel audio data simultaneously to support real-time communication with various natural turn-taking events such as interruptions Chen et al. (2025), backchannels Défossez et al. (2024), and overlapping speech Nguyen et al. (2023), aiming to emulate human-like conversations Nguyen et al. (2023); Wang et al. (2025). Existing FD-SLMs can be categorized as cascaded and end-to-end (e2e) systems. Cascaded FD-SLMs utilize a separate prediction module to classify spoken dialogues into predefined states, guiding SLM generation Chen et al. (2025); Wang et al. (2024). While demonstrating efficacy in handling elementary turn-taking like interruptions, they are inherently bound by their dependence on explicit rule-based frameworks. In contrast, e2e FD-SLMs, which we adopt, leverage real-world conversational data to train models e2e, capturing nuanced dialogue patterns for more natural, human-like interactions Wang et al. (2025); Veluri et al. (2024). Figure 1 illustrates these two FD-SLM types.
While e2e FD-SLMs enable more natural conversational experiences, a key challenge is that their conversational abilities degrade compared to text-based Large Language Models (TLMs), due to prolonged speech sequences and the scarcity of high-quality speech training data Zeng et al. (2024); Xie and Wu (2024). One promising solution is to transfer text abilities to speech abilities via text-speech interleaved generation, where the model first generates the assistant’s text response and uses it to guide subsequent speech generation. However, applying text guidance in e2e FD-SLMs poses significant challenges, as these models require precise time alignment between user and assistant audio streams to capture complex conversational patterns and provide timely, meaningful responses. Interleaving text tokens into the audio streams can easily disrupt this alignment, compromising the responsiveness and coherence of the generated speech.
The challenges of integrating text into assistant-channel audio data can be summarized in two aspects. 1) Insertion timing. Text must be inserted precisely when the assistant begins speaking. If it is inserted too early, the assistant may respond without the full user context, causing hallucinations. If it is inserted too late, the assistant’s speech may precede the text, undermining the text guidance’s effectiveness. 2) Insertion length. The length of the inserted text is also critical. If it is too long, the assistant may overspeak and fail to adapt to subsequent user inputs; if too short, the semantic content becomes fragmented, weakening the effectiveness of text guidance Yao et al. (2025).
To address these challenges, we propose a novel approach, TurnGuide, to seamlessly integrate text data into real-world double-channel spoken conversational data for the training of e2e FD-SLMs. Specifically, we dynamically segment the assistant’s speech into turns, which represent the reasonable lengths of spoken content that convey the assistant’s intended expression. Then, the assistant is trained to interleave the text-speech generation for each turn for proper guidance. By doing so, our method effectively addresses both insertion timing and insertion length challenges. Extensive experiments show that our approach significantly improves the conversational abilities of e2e FD-SLMs by not only producing semantically meaningful and coherent speech but also achieving state-of-the-art performance on various turn-taking events. To summarize our contributions:
  • We propose TurnGuide, a novel text-speech interleaved generation approach for e2e FD-SLMs grounded in natural speech data.
  • We introduce a dynamic turn segmentation framework and a text-guided dialogue modeling framework within TurnGuide to address the insertion timing and length in double-channel dialogues.
  • We demonstrate through extensive experiments that TurnGuide substantially improves not only FD-SLMs’ ability to produce semantically meaningful, coherent speech but also the behaviors on various turn-taking events.
2 Related Works
2.1 Full Duplex Speech Language Models
FD-SLMs process double-channel conversational data to interact with users seamlessly and can be categorized into cascaded and end-to-end (e2e) approaches.
2.1.1 Cascaded FD-SLMs
Cascaded FD-SLMs utilize a separate prediction module to classify full-duplex dialogue into a finite number of predefined states, which are typically designed to facilitate transitions between listening and speaking to guide SLM behavior. Training cascaded FD-SLMs consists of two key components: a state predictor, which predicts the SLM’s next state, and a backbone SLM, typically trained on synthetic conversational data. Therefore, cascaded FD-SLMs are constrained by their predefined behaviors, preventing them from authentically engaging in human-like spoken conversations.
The main focus of cascaded FD-SLMs is on defining states. For instance, LSLM Ma et al. (2025) uses states like “interrupted” and “keep speaking”, where interruptions are triggered by specific keywords. Mini-Omni 2 Xie and Wu (2024) specifies “start speaking” and “stop speaking”, halting generation upon hearing “Stop Omni!”. Freeze-Omni Wang et al. (2024) bases states on whether the user has finished speaking, prompting the model to wait or interrupt. Minmo Chen et al. (2025) simplifies its states to whether the model should speak up or keep listening. All these methods rely on synthetic conversational data to manage state transitions, limiting their ability to foster natural, human-like spoken interactions.
2.1.2 End-to-end FD-SLMs
E2e FD-SLMs, on the other hand, employ a transformer-based foundation model architecture Vaswani et al. (2017); Achiam et al. (2023) to learn from real-world double-channel conversational audio data. This e2e approach enables SLMs to autonomously capture intricate conversational dynamics from authentic human interactions, enabling more natural conversations.
The main focus of e2e FD-SLMs is on effectively modeling double-channel audio data and capturing the interactions between these channels. dGSLM Nguyen et al. (2023) introduces a Dialogue Language Model that employs distinct transformers to process each audio channel separately, using cross-attention mechanisms to model their interactions. SyncLLM Veluri et al. (2024) takes a different approach by jointly modeling the double-channel audio within a single LLM, interleaving the two audio sequences into a unified sequence. Moshi Défossez et al. (2024) adopts an RQ-Transformer architecture Lee et al. (2022), designed to predict the assistant channel’s output based on the user channel’s input at each time step. NTPP Wang et al. (2025) proposes a Next-Token-Pair Prediction training objective, enabling simultaneous prediction of audio tokens for both channels in each step.
2.2 Text-Guided Speech Generation in SLMs
SLMs enable e2e spoken interactions by generating output speech based on user speech input. However, studies have shown that SLMs exhibit significantly lower performance compared to TLMs counterparts Cui et al. (2025); Chen et al. (2025); Long et al. (2025). To mitigate this, researchers focus on integrating text to guide the speech generation process, aiming to enhance SLMs by transferring capabilities from TLMs.
Transferring abilities from TLMs to SLMs involves two key processes: representation alignment and text-speech interleaved generation. Representation alignment ensures the connection between text and speech representations within the SLM backbone, enabling knowledge transfer from text modality to speech modality. This alignment is commonly achieved through Automatic Speech Recognition (ASR) and Text-To-Speech (TTS) tasks during SLM pre-training Zeng et al. (2024); Wang et al. (2024); Chen et al. (2025). Some approaches also partially replace speech data with its corresponding text in the pre-training data Nguyen et al. (2025); Zeng et al. (2024). Subsequently, SLMs are trained to answer spoken queries by interleaving the generation of textual output with speech output, thereby effectively leveraging text information to enhance the quality of spoken responses (Zeng et al., 2024; Défossez et al., 2024).
The exploration of the text-guided speech generation for e2e FD-SLMs remains in its infancy. Currently, the only notable study in this area is Moshi Défossez et al. (2024), which proposes adding text tokens corresponding to their respective speech sounds as they occur. However, this method fragments text semantics excessively and does not significantly improve SLM performance. In contrast, our approach aggregates the text semantics and inserts text by turns, achieving substantial enhancements in SLM performance.
3 Methodology
Figure 2: Illustration of the TurnGuide approach. The first part illustrates the multi-modal turn segmentation and alignment framework, which dynamically segments assistant speech into turns and aligns the text with speech turns. The second part shows the text-guided full duplex dialogue modeling framework with two interleaving strategies.
We present the comprehensive methodology of TurnGuide, designed to enhance e2e FD-SLMs by utilizing text guidance. This approach is divided into two key frameworks: one for dynamically segmenting spoken dialogue into turns and performing speech-text data alignment, and the other for modeling text-guided full-duplex dialogue.
3.1 Dynamic Turn Segmentation and Alignment Framework
To enable text guidance in full-duplex speech interactions, it is essential to derive training data that effectively combines text and speech. As a first step, it is crucial to determine how text can be incorporated without disrupting the natural conversational flow. This requires the text to be introduced at an appropriate moment and kept within a reasonable length. We propose tackling these challenges by segmenting the speech into turns and aligning each part of the text with its corresponding speech turn.
As shown in the first part of Figure 2, the framework starts by applying Voice Activity Detection (VAD) to identify speech segments within the audio stream. Let $A$ represent the input audio sequence. VAD produces a list of speech segments:
$V=\text{VAD}(A)=\{(\tau_{1}^{start},\tau_{1}^{end}),(\tau_{2}^{start},\tau_{2}^{end}),$...\}, (1)
where each tuple $(\tau_{i}^{start},\tau_{i}^{end})$ represents the start and end timestamps of the $i$-th detected speech segment. We use the pyannote library Bredin et al. (2020) as the VAD module. Considering the VAD result is too fragmented, we merge nearby speech segments into Inter-Pausal Units (IPUs) Nguyen et al. (2023), which are defined by consecutive segments having pauses within a threshold $\tau_{IPU}=0.5$ seconds, i.e.,
$IPU_{j}=\bigcup_{i\in G_{j}}(\tau_{i}^{start},\tau_{i}^{end}),$ (2)
where $G_{j}=\{i:\tau_{i+1}^{start}-\tau_{i}^{end}<\tau_{IPU}\}$ defines the group of consecutive segments to be merged into IPU $j$.
Enabling real-time text guidance requires the precise alignment of text and speech data. We use an Automatic Speech Recognition (ASR) module with word-level timestamp functionality to acquire the time-aligned transcript. While some speech dialogue datasets already provide time-aligned text data, our approach ensures adaptability to unlabeled dialogue datasets. The ASR process is formulated as
$W=\text{ASR}(A)=\{w_{1},w_{2},$...\}, (3)
where $w_{i}$ represents the $i$-th word in the transcription. The word-level timestamp extraction process is formulated as
$\Theta=\text{TE}(A)=\{(\delta_{1}^{start},\delta_{1}^{end}),(\delta_{2}^{start},\delta_{2}^{end}),$...\}, (4)
where $(\delta_{i}^{start},\delta_{i}^{end})$ represents the start and end timestamps of word $w_{i}$. In practice, we employ the Whisper medium model Radford et al. (2023) for both ASR and timestamp extraction tasks. To enhance the stability of the timestamp extraction process, we utilize the whisper-timestamped package Louradour (2023).
Finally, we align and group each word into its associated IPU using $\tau^{start}$. Additionally, we incorporate a tolerance parameter $\tau_{tol}=0.6$ seconds to adjust the word segmentation boundary slightly. This adjustment accounts for instances where, in practice, the first word in an IPU might incorrectly align with the preceding IPU due to inconsistencies in timestamping. The process is defined as
$W_{j}=\{w_{i}:\tau_{IPU_{j}}^{start}-\tau_{tol}\leq\delta_{i}^{start}\leq\tau_{IPU_{j+1}}^{start}-\tau_{tol}\},$ (5)
where $W_{j}$ represents words grouped into turn $j$, and $\tau_{IPU_{j}}^{start}$ and $\tau_{IPU_{j+1}}^{start}$ represent the start times of IPU $j$ and $j+1$. We denote the speech segment corresponding to IPU $j$ as $S_{j}$. The final output consists of aligned text-speech pairs $\{(W_{j},S_{j})\}$ where each text segment is temporally aligned with its corresponding speech segment, serving as training data for text-guided full-duplex dialogue modeling.
3.2 Text-Guided Full Duplex Dialogue Modeling Framework
After deriving the aligned dialogue turns, the next step is to model double-channel audio data with text guidance. We use GLM-4-Voice Zeng et al. (2024), an off-the-shelf SLM, as the backbone for training on full-duplex dialogue datasets. GLM-4-Voice includes a speech tokenizer to convert audio into discrete tokens, a language model to process these tokens, and a vocoder to reconstruct speech. Its strong speech-text alignment makes it ideal for the backbone of text-guided speech interaction. As illustrated in the second part of Figure 2, our text-guided full duplex dialogue modeling framework centers around two interleaving strategies:
Channel-wise Interleaving. Full-duplex dialogue modeling necessitates the simultaneous processing of two-channel audio to capture cross-channel interactions. A common technique to address this is by interleaving the two audio channels into a single sequence, either with token-level interleaving Wang et al. (2025) or chunk-level interleaving Veluri et al. (2024); Zhang et al. (2024). We choose chunk-level over token-level interleaving because it demonstrates better performance in our experiments. We first consider channel-wise interleaving for speech tokens. To implement this technique, let $s_{u}$ and $s_{a}$ represent speech tokens from the user and assistant channels. Each channel’s speech tokens are divided into speech chunks with a chunk size of $\tau_{SC}=5$. Given GLM-4-Voice’s frame rate of $fr=12.5\text{Hz}$, each speech chunk corresponds to a 400-ms audio segment. The chunk-interleaving process alternates segments between the two channels, forming a sequence $SC=[SC_{u}^{1},SC_{a}^{1},SC_{u}^{2},SC_{a}^{2},\dots]$. Each speech chunk contains $\tau_{SC}$ speech tokens, with a unique speaker ID token inserted at the beginning to distinguish between the channels. Specifically, $SC_{u}^{i}=[UID,s_{u}^{i\times\tau_{SC}+1},s_{u}^{i\times\tau_{SC}+2},\dots,s_{u}^{i\times\tau_{SC}+\tau_{SC}}]$ and $SC_{a}^{i}=[AID,s_{a}^{i\times\tau_{SC}+1},s_{a}^{i\times\tau_{SC}+2},\dots,s_{a}^{i\times\tau_{SC}+\tau_{SC}}]$ represent the $i$-th user and assistant channel speech chunks, where $UID$ and $AID$ denote the user and assistant ID tokens.
Text-speech Interleaving. We apply text guidance on the assistant channel. Specifically, we interleave the text tokens with speech tokens for each turn, starting from the first speech chunk of the turn.
Given the text in a turn $W_{j}$, we first split it into text chunks $TC_{j}=[TC_{j}^{1},TC_{j}^{2},$...] with chunk size $\tau_{TC}=5$, where $TC_{j}^{k}=[t_{j}^{k\times\tau_{TC}+1},t_{j}^{k\times\tau_{TC}+2},...,t_{j}^{k\times\tau_{TC}+\tau_{TC}}]$ represents the $i$-th text chunk for turn $j$. Then, we append an “end of turn” ($\langle EOT\rangle$) token to the last text chunk, and append an “end of chunk” ($\langle EOC\rangle$) token to the rest of the text chunks. This is to force the model to learn the turn boundary information. Second, we calculate the corresponding speech chunk for text-speech interleaving for each text chunk $TC_{j}^{k}$ in $W_{j}$ as
$k_{j}=\lfloor\frac{\tau_{j}^{k0}\times fr}{\tau_{SC}}\rfloor,$ (6)
where $\tau_{j}^{k0}$ represents the timestamp for the first text token of the $k$-th text chunk in turn $j$. The text-speech interleaving is performed at the $k_{j}$-th chunk of every turn $j$.
The final sequence is formed by chunks with the two interleaving strategies, $FC=[C_{u}^{1},C_{a}^{1},C_{u}^{2},C_{a}^{2},$...], where
$C_{u}^{i}=SC_{u}^{i}\text{ and }C_{a}^{i}=\begin{cases}[TC_{j}^{k},SC_{a}^{i}]&\text{if text chunk presents}\\ SC_{a}^{i}&\text{otherwise}.\end{cases}$
The decision to use a 5:5 ratio for text chunks and speech chunks ensures that text generation precedes speech generation, providing effective text guidance. In summary, our full duplex dialogue modeling framework leverages two interleaving strategies to enable both cross-channel interaction and text-guided speech modeling.
SLMs Fisher-finetuned Open-source
dGSLM Nguyen et al. (2023)
SyncLLM Veluri et al. (2024)
Moshi Défossez et al. (2024)
OmniFlatten Zhang et al. (2024)
NTPP Wang et al. (2025)
Wu et al. Wu et al. (2025)
Table 1: Summary of the open-source status of previous e2e FD-SLMs papers.
SLM Unconditional Conditional - All Conditional - Assistant Overall
0.8 0.9 1.0 1.3 AVG 0.8 0.9 1.0 1.3 AVG 0.8 0.9 1.0 1.3 AVG
dGSLM 4.70 5.00 4.94 4.03 4.67 - - - - - - - - - - -
STI 4.06 4.94 6.00 7.79 5.70 3.58 4.36 4.94 6.18 4.76 2.76 3.36 3.76 5.24 3.79 4.76
SCI 5.48 6.58 7.52 8.21 6.94 5.15 5.67 6.00 6.55 5.85 4.09 4.61 5.06 6.06 4.94 5.91
Moshi - - - - - 6.70 6.39 5.94 4.21 5.82 6.39 6.42 6.30 4.24 5.85 -
Moshi TS 5.21 6.24 7.30 8.30 6.76 4.42 4.94 5.52 6.45 5.33 3.73 4.18 4.76 5.94 4.67 5.58
TurnGuide 8.12 8.76 8.88 8.61 8.61 6.24 6.58 7.09 7.15 6.76 6.12 6.42 6.58 6.79 6.48 7.27
TurnGuide L2:1 8.70 9.27 9.33 8.76 9.03 6.52 7.03 7.30 7.21 7.03 6.61 7.06 7.30 7.12 7.03 7.70
TurnGuide L3:1 8.91 9.24 9.33 8.67 9.03 6.97 7.24 7.30 7.24 7.18 6.88 7.21 7.21 7.18 7.12 7.79
TurnGuide GT Text 7.91 8.39 8.82 8.55 8.42 6.73 7.03 7.15 6.97 6.97 6.36 6.94 6.97 6.97 6.82 7.39
Table 2: Semantic evaluation results (0-10↑) of different FD-SLMs under different temperature settings by GPT-4o.
4 Experiments
In this section, we assess the effectiveness of TurnGuide, focusing on the semantic meaningfulness and the turn-taking behaviors of the generated dialogues.
4.1 Experimental Setups
4.1.1 Training Data
We utilize the Fisher dataset Cieri et al. (2004), a double-channel telephone conversation corpus containing 2,000 hours of audio data, to train TurnGuide. Following prior works (Nguyen et al., 2023; Veluri et al., 2024), we pre-process each recording in the Fisher dataset into segments of 120-second clips for training and evaluation. We divide the Fisher dataset into training, validation, and test sets, with detailed information provided in Appendix E.
4.1.2 Baselines and Our Approach
We consider the following models or methods as our baselines. 1) dGSLM Nguyen et al. (2023): a dual-Transformer architecture, each modeling an individual audio channel. 2) Speech Token Interleaving (STI): We train the GLM-4-Voice backbone on the Fisher dataset by interleaving the speech tokens from the two channels, one token at a time. 3) Speech Chunk Interleaving (SCI): Similar to Speech Token Interleaving, but it interleaves speech tokens chunk by chunk with a chunk size of 5. 4) Text-speech Token Alignment (Moshi Training Strategy, Moshi TS Défossez et al. (2024)): The only existing method for text-guided speech generation in e2e FD-SLMs. It involves inserting each text token exactly at the time its corresponding sound is produced. However, Moshi is trained using a wide range of datasets, with Fisher being only one of them, making it not directly comparable to our Fisher-finetuned baselines. To address this, we train the GLM-4-Voice backbone on the Fisher dataset with Moshi’s training methodology to compare it with other baselines. Our detailed implementation of Moshi TS is illustrated in Appendix B. 5) Moshi: We also include the original Moshi model for completeness. 6) Our Approach (TurnGuide): We train the GLM-4-Voice backbone on the Fisher dataset by utilizing the text-guided full duplex dialogue modeling framework introduced in Section 3.2. The text-speech dataset is pre-processed using the framework illustrated in Section 3.1.
Remark 1
While there are other e2e FD-SLMs, most have not been (fully) open-sourced yet, as summarized in Table 1. Therefore, we mainly evaluate our approach against dGSLM, Moshi, and our trained baselines.
4.1.3 Model Training and Inference
The models are trained on the Fisher dataset for 2 epochs with a global batch size of 256, a learning rate of 4e-6, a cosine learning rate scheduler, no weight decay, and 20 warmup steps. During inference, we sample 1,000 audio clips from the Fisher test sets. The first 30 seconds serve as prompts, while the model generates 90-second continuations using two methods: unconditional and conditional generation. In unconditional generation, we train the model with text guidance in both audio channels, and the model generates both channels for the 90-second continuation during inference. This setting has potential in two-speaker podcast generation. In conditional generation, we train the model with only assistant-channel text guidance, and the model generates the assistant-channel audio based on the user-channel audio as context, which can be used for spoken interaction purposes.
4.2 Results
We aim to answer three critical Research Questions (RQs) through comprehensive experiments. RQ1: Does our method significantly improve the semantic meaningfulness and coherence of the generated full-duplex dialogue? RQ2: Can our method preserve the natural conversational flow after integrating additional texts into the sequence? RQ3: How does the model perform if we insert text data with different timings and lengths?
Table 3: Full-Duplex-Bench evaluation results. The best performance is bold, and the second-best performance is underlined. Latency is presented in seconds. The scores in brackets indicate temperature.
Dimension Pause Handling Backchannel Smooth Turn Taking User Interruption
Data Candor ICC Candor Synthetic
Metric TOR ($\downarrow$) TOR ($\downarrow$) Freq ($\uparrow$) JSD ($\downarrow$) TOR ($\uparrow$) Latency ($\downarrow$) TOR ($\uparrow$) GPT-4o($\uparrow$) Latency ($\downarrow$)
dGSLM 0.935 0.691 0.015 0.934 0.975 0.352 0.917 0.201 2.531
Moshi 0.980 1.000 0.001 0.957 0.941 0.265 1.000 0.765 0.257
TurnGuide L2:1 (0.8) 0.390 0.667 0.013 0.927 0.773 0.567 0.558 1.349 2.263
TurnGuide L2:1 (0.9) 0.441 0.519 0.019 0.860 0.773 0.357 0.689 1.441 1.870
TurnGuide L2:1 (1.0) 0.587 0.630 0.021 0.887 0.807 0.282 0.866 1.471 1.566
TurnGuide L2:1 (1.3) 0.789 0.500 0.043 0.814 0.950 0.102 1.000 1.150 0.900
Figure 3: Perplexity evaluation of FD-SLMs across temperatures using Fisher-text-only, GLM-4-Voice, and Llama-3.1-8B-Instruct as evaluation models.
4.2.1 RQ1
To evaluate the semantics of the generated dialogue, the most intuitive method is to conduct human evaluations. However, considering the time, cost, and potential inconsistencies associated with human evaluations, we propose utilizing the GPT-score. Specifically, given a generated double-channel dialogue audio, we first transcribe the audio from each channel into text with sentence-level timestamps. For this process, we employ the stable-ts package Jian (2023) with the Whisper-large-v3 model. Subsequently, the transcribed sentences from both channels are grouped based on their start times, effectively representing the dialogue between the two speakers in text form along the timeline. Finally, the consolidated text is analyzed using GPT-4o Hurst et al. (2024), which assesses the meaningfulness and coherence of the dialogue on a scale of 0 to 5 with detailed rubrics. The complete GPT-4o prompts for unconditional and conditional generation settings are shown in Appendix G. To demonstrate the reliability of GPT-based evaluation, we conduct human evaluations, showing strong consistency between human- and GPT-evaluated scores (see Appendix A).
We demonstrate the effectiveness of TurnGuide through evaluations in both unconditional and conditional settings. In the “Conditional - All” scenario, GPT-4o uses ordered transcripts from both channels to assess responses based on user speech. However, we observe that GPT-4o overly prioritizes user-assistant interactions, neglecting the coherence of assistant-generated speech within its channel. To address this, the “Conditional - Assistant” setting provides only the assistant-channel transcript, enabling a focused analysis of semantic quality within the assistant channel. We also observe that GPT-4o tends to assign conservative scores for this task, which is evidenced by the ground truth text from the Fisher dataset achieving an average GPT-score of 3.3. Therefore, we rescale all GPT-score results from 0-3.3 to 0-10.0.
We compare TurnGuide with various baselines under different temperatures (as shown in Table 2) and make the following observations. 1) Significant performance improvements over baselines. TurnGuide performs significantly better than SCI, Moshi TS (over 30% performance gain) and Moshi (over 20% performance gain), showing that TurnGuide successfully leverages the text information to boost the meaningfulness and coherence of the generated dialogue. 2) Speech-only methods and text-speech methods excel at different temperatures. SCI consistently performs the best at the largest temperature (1.3), whereas TurnGuide tend to perform the best at temperature 1.0. Additionally, TurnGuide shows more performance gain under lower temperatures, such as 0.8 and 0.9. 3) Focusing more on text generation makes the model perform better. In “TurnGuide L2:1" and “TurnGuide L2:1" settings, we set the training loss weight of text and speech tokens to be 2:1 and 3:1, and the results show that they consistently outperform the standard 1:1 setting (“TurnGuide"). 4) TurnGuide achieves performance comparable to its upper bounds. In Table 2, “TurnGuide GT Text" refers to the GLM-4-Voice model trained with our proposed framework but switches to the ground truth transcriptions from the Fisher dataset. TurnGuide performs similarly with TurnGuide GT Text in the 1:1 loss setting (7.27 vs. 7.39), and the TurnGuide 2:1 and 3:1 outperform the upper bound in 1:1 loss.
Further validation with perplexity. To facilitate the semantic evaluation using GPT-score, perplexity is employed to further demonstrate the superior semantic abilities of TurnGuide. Specifically, we calculate perplexity using the same ordered transcript used in the GPT-score evaluation. Perplexity is measured with three distinct language models: Fisher-text-only, GLM-4-Voice, and Llama-3.1-8b-Instruct Dubey et al. (2024). The Fisher-text-only model, trained using Fisher-provided text transcript data on the GLM-4-Voice base model, serves as the ground truth for text-based dialogue distribution. GLM-4-Voice acts as the SLM backbone, while Llama-3.1-8b-Instruct serves as the text-only model backbone. We analyze the perplexity of generated dialogues across different temperature settings in both conditional and unconditional scenarios. As illustrated in Figure 3, TurnGuide consistently achieves lower perplexity compared to baseline approaches, demonstrating superior semantic modeling abilities.
4.2.2 RQ2
Evaluating the naturalness of FD-SLM interaction involves assessing various turn-taking events in the full-duplex dialogues they generate. Traditional FD-SLM studies (Nguyen et al., 2023; Veluri et al., 2024; Défossez et al., 2024) often analyze turn-taking behaviors at the corpus level, comparing the model-generated statistics to the corresponding ground-truth statistics. However, Lin et al. (2025) highlights that corpus-level statistics are often difficult to interpret, making them less suitable as an evaluation method. To overcome this limitation, they introduce the Full-Duplex-Bench benchmark, designed to evaluate the fine-grained turn-taking behaviors of FD-SLMs. This benchmark specifically assesses core turn-taking abilities, such as Smooth Turn-taking, User Interruption, Pause Handling, and Backchanneling. The detailed explanation of the four turn-taking events is in Appendix H.
We compare TurnGuide trained with 2:1 loss with the baseline approaches on Full-Duplex-Bench (as shown in 3) and highlight the following. 1) TurnGuide demonstrates superior interaction naturalness compared to baseline approaches. TurnGuide surpasses baselines on nearly all the metrics of the four turn-taking events, except for the TOR of smooth turn-taking and the latency of user interruption. 2) TurnGuide achieves optimal turn-taking performance when the temperature is set to 1.3. At this temperature setting, TurnGuide outperforms not only the baselines but also TurnGuide at other temperatures.
SLM Conditional - All
0.8 0.9 1.0 1.3 AVG
TurnGuide 6.24 6.58 7.09 7.15 6.76
Too Long 5.73 5.97 6.30 6.52 6.12
Too Short 6.06 6.52 6.73 6.70 6.52
Too Early 4.88 5.06 5.64 6.88 5.61
Too Late 5.27 5.76 6.06 6.70 5.94
Table 4: Semantic evaluation results of TurnGuide on different text insertion timings and lengths.
4.2.3 RQ3
We aim to support the claim that the text insertion in the full duplex dialogue modeling process exhibits the insertion timing and length challenges. Specifically, we simulate the following four text insertion scenarios and demonstrate the superiority of our approach. 1) Too Long: Each text chunk contains 20 text tokens. 2) Too Short: Each text chunk contains 2 text tokens. 3) Too Early: The insertion time of every text chunk is shifted to 4s earlier. 4) Too Late: The insertion time of every text chunk is shifted to 4s later.
The results presented in Table 4 highlight the following. 1) Text insertion timings and lengths indeed pose challenges to the full duplex dialogue modeling process. We observe that all the simulated settings significantly degrade the model performance. 2) Model performance is particularly sensitive to insertion timing. The Too Early and Too Late scenarios result in greater performance degradation compared to Too Long and Too Short.
5 Conclusion
We present TurnGuide, a novel method to improve the dialogue generation abilities of e2e FD-SLMs. By segmenting assistant speech into turns and training the model with channel-wise interleaving and turn-level text-speech interleaving strategies, our approach effectively addresses key challenges of text insertion timing and length, ensuring coherent and natural dialogue generation. Experiments show significant improvements of our method in semantic quality and various natural turn-taking events, validating its effectiveness for spoken dialogue systems and paving the way for more meaningful and human-like speech interactions.
Limitations
While the paper introduces TurnGuide, a novel dynamic turn-level text-speech interleaved generation approach for enhancing e2e FD-SLMs, it still presents several limitations. First, the effectiveness of TurnGuide is primarily validated on the Fisher dataset, which, although widely used, may not fully capture the diversity and complexity of real-world conversational scenarios since Fisher is a telephone conversation dataset. However, addressing this limitation would necessitate the effort of open-sourcing real-world double-channel dialogue datasets that encompass a broader range of styles. Second, the evaluation relies heavily on GPT-4o-based automated metrics (GPT-score) for semantic assessment, which, despite being aligned with human judgments in most cases, may not perfectly reflect the nuanced human perceptions of dialogue quality.
We then suggest several promising directions to expand upon TurnGuide. One key avenue is to explore the generalization of the proposed framework to more diverse and challenging conversational datasets, including those featuring multiple speakers, varied languages, or more spontaneous dialogue styles. Enhancing robustness to noisy or low-resource settings—such as improving alignment without high-quality transcriptions or adapting to imperfect voice activity detection—would also broaden real-world applicability. Moreover, integrating additional modalities (e.g., visual cues or gestures) could further enrich dialogue modeling and understanding.
Appendix A Judgement Alignment between Human and GPT-4o in Semantic Evaluation
Difference 1 2 3 Overall
Accuracy 7/10 9/10 9/10 25/30
Table 5: Alignment accuracy between human judgments and GPT-4o semantic evaluation scores for generated dialogue pairs with varying score differences.
SLMs M-MOS N-MOS
SCI 2.24 2.47
Moshi 2.69 2.75
TurnGuide 3.26 3.23
Ground Truth 4.25 4.29
Table 6: Absolute rating results of the human evaluations assessing the Meaningfulness (M-MOS) and Naturalness (N-MOS) of the dialogues.
Figure 5: Training and validation loss comparison of different training strategies.
Figure 6: Illustration of the four turn-taking events used in the evaluation.
In this paper, we primarily utilize GPT-4o to evaluate the semantics of the generated dialogues. This raises a crucial question: Can AI evaluation scores accurately capture human preferences? To investigate this, we conduct two human listening tests, including pairwise comparison and absolute rating, to examine the alignment between GPT-evaluated scores and human judgments.
A.1 Pairwise Comparison
Pairwise comparison involves organizing the generated dialogues into pairs and having human judges assess which dialogue in each pair exhibits better semantic quality. This experiment aims to determine whether GPT-4o can effectively differentiate between good and bad dialogues. Specifically, we randomly select 30 pairs of generated dialogues that vary in their GPT-scores: 10 pairs with a GPT-score difference of 1, 10 pairs with a difference of 2, and 10 pairs with a difference of 3. To ensure unbiased evaluation, the order of dialogues within each pair is randomized. Additionally, judges are provided with both the audio of the dialogues and their transcripts, generated using Whisper-large-v3 with the stable-ts package. We recruit three human judges, all of whom have native English proficiency, to carry out the evaluations. All human raters provided informed consent before participating and were informed that their annotations may be used for academic research and reported in an academic publication. The complete instructions given to the participants are presented in Figure 9.
The results highlight strong human-AI alignment in this task, as presented in Table 5. Notably, human judgments match GPT-4o in 25 out of the 30 pairs evaluated. Among the mismatches, most occur in cases with minor differences (1-difference), but the judgments still match for 7 out of 10 pairs. This suggests GPT excels at distinguishing dialogues with significant semantic differences, but it may struggle when dialogues exhibit similar semantic levels. These findings align closely with human evaluations, underscoring the reliability of our GPT-score evaluation methodology.
A.2 Absolute Rating
Absolute rating involves human evaluators directly assessing the quality of dialogues generated by various models using a 5-point Likert scale. In this method, we can gather the human-evaluated scores and compare them with the GPT-scores in Table 2. A strong alignment between humans and AI is observed when the ranking of scores for each model remains consistent across both human and AI evaluations. We randomly choose 10 audio prompts and obtain their continuations by SCI, Moshi, TurnGuide, and the Ground Truth, under each model’s best-performing temperature. This results in a total of 40 audio samples. Following the exact settings and instructions outlined in Veluri et al. (2024), human evaluators rate the Meaningfulness (M-MOS) and Naturalness (N-MOS) of the dialogues. Same as pairwise comparison, judges are provided with both the audio of the dialogues and their transcripts. We recruit nine human judges with native English proficiency for this task.
The results presented in Table 6 also demonstrate a strong alignment between human and AI evaluations. Notably, the ranking of the models aligns with the AI’s assessments in Table 2, further validating the reliability of our automated semantic evaluation approach. Moreover, among all the models, TurnGuide achieves the highest scores in both meaningfulness and naturalness, underscoring its effectiveness as a superior general-purpose FD-SLM.
Appendix B Moshi TS Implementation
As detailed in Section 4.1, we do not directly compare our model with Moshi, as Moshi is a general-purpose SLM that utilizes different pre-training speech data and is not fine-tuned on the Fisher dataset during its final training stage. Instead, we opt to train a model using Moshi’s training strategy, employing GLM-4-Voice on the Fisher dataset. This approach ensures a fair comparison, as the base model and training dataset remain identical, with the only difference between our approach and Moshi TS being the training strategy used. Therefore, in this section, we illustrate the details of the training process of Moshi TS.
First, Moshi utilizes eight codebooks, where every time step includes eight speech tokens for each channel. We adapt this technique for the GLM-4-Voice backbone, which relies on a single codebook method, allotting one speech token per channel per time step. Then, the tokens from each channel are interleaved at each time step, making Moshi rely on a token-level interleaving approach. In the original implementation of Moshi, a text-channel token precedes each assistant-channel speech token, which could either be a genuine text token, a text pad token, or an end-of-text pad token. Moshi also employs text-speech interleaving at token-level, where each text token is incorporated at the precise moment it is spoken. Consequently, if a time step features an audible text token, the corresponding text token is inserted. If not, either a text pad token or an end-of-text pad token is inserted, determined by whether the subsequent text-channel token is an actual text token. We follow the original Moshi implementation by extracting the timestamp of each word with Whisper-timestamped package Louradour (2023) and the medium Whisper model Radford et al. (2023). The illustration of our implemented Moshi TS is shown in Figure 4.
SLM Occurrence Cumulative Duration Averaged Duration Overall
0.8 0.9 1.0 1.3 AVG 0.8 0.9 1.0 1.3 AVG 0.8 0.9 1.0 1.3 AVG
Unconditional
dGSLM 0.33 0.41 0.45 0.36 0.39 0.45 0.53 0.55 0.40 0.48 0.30 0.37 0.42 0.26 0.34 0.40
STI 0.40 0.47 0.43 0.46 0.44 0.39 0.37 0.38 0.48 0.41 0.25 0.22 0.23 0.27 0.24 0.36
SCI 0.57 0.57 0.61 0.62 0.59 0.48 0.48 0.52 0.58 0.51 0.26 0.25 0.33 0.36 0.30 0.47
Moshi TS 0.52 0.53 0.56 0.61 0.56 0.38 0.43 0.46 0.54 0.45 0.20 0.24 0.28 0.34 0.26 0.42
TurnGuide 0.57 0.57 0.63 0.62 0.60 0.46 0.52 0.53 0.56 0.52 0.27 0.25 0.34 0.36 0.30 0.47
TurnGuide L2:1 0.57 0.60 0.62 0.62 0.60 0.50 0.51 0.55 0.56 0.53 0.29 0.26 0.36 0.37 0.32 0.48
TurnGuide L3:1 0.58 0.59 0.63 0.58 0.60 0.47 0.50 0.54 0.55 0.51 0.31 0.25 0.33 0.37 0.32 0.47
Conditional
STI 0.50 0.52 0.58 0.62 0.56 0.40 0.42 0.47 0.53 0.46 0.20 0.21 0.27 0.32 0.25 0.42
SCI 0.47 0.49 0.49 0.50 0.49 0.28 0.40 0.46 0.53 0.42 0.17 0.25 0.29 0.34 0.26 0.39
Moshi 0.65 0.62 0.55 0.51 0.58 0.43 0.43 0.46 0.35 0.42 0.32 0.29 0.35 0.33 0.32 0.44
Moshi TS 0.48 0.48 0.48 0.45 0.47 0.39 0.39 0.43 0.50 0.43 0.24 0.25 0.29 0.29 0.27 0.39
TurnGuide 0.46 0.50 0.51 0.46 0.48 0.38 0.46 0.48 0.50 0.45 0.24 0.26 0.34 0.32 0.29 0.41
TurnGuide L2:1 0.50 0.48 0.51 0.43 0.48 0.39 0.49 0.54 0.49 0.48 0.27 0.33 0.33 0.33 0.31 0.42
TurnGuide L3:1 0.46 0.48 0.49 0.47 0.48 0.39 0.46 0.51 0.50 0.46 0.20 0.30 0.28 0.32 0.27 0.41
Table 7: Pearson correlation results of corpus-level turn-taking events with unconditional and conditional generation of various FD-SLMs evaluated under different temperature configurations.
Appendix C Comparison of Different Training Strategies
In this section, we provide a comparison of different training strategies and share corresponding insights. We start by examining the impact of these strategies on the training process by analyzing the training loss curves displayed in Figure 5. First, SCI achieves a significant reduction in training loss compared to STI. This improvement can likely be attributed to the addition of speaker ID tokens for each channel, which provides extra information about the correspondence between speech tokens and their respective speakers, thereby simplifying the model’s learning process. Furthermore, the reduction in training loss may also result from the organization of speech tokens into chunks, which enhances contextual coherence compared to STI. Second, TurnGuide demonstrates even lower training loss than STI. More specifically, TurnGuide outperforms both SCI and STI, with configurations such as TurnGuide L2:1 and TurnGuide L3:1 achieving even greater reductions in training loss. These results highlight the benefits of the text-speech alignment process, which facilitates SLM training and enables superior semantic performance. Finally, the loss reduction observed in Moshi TS does not translate into performance improvements. While Moshi TS achieves one of the lowest training losses, its overall performance trails behind our approach. This discrepancy may stem from the excessive presence of text pad tokens relative to actual text tokens. The repetitive nature of pad tokens simplifies prediction, leading to artificially low training losses without corresponding gains in real-world performance.
We then intuitively explore the reasons behind the lower performance of Moshi TS compared to TurnGuide and identify three key factors contributing to this difference. First, Moshi TS tends to fragment text semantics. Both Moshi TS and our method aim to enhance the semantics of generated speech using textual information. However, Moshi TS splits text tokens so significantly that it hampers the model’s ability to learn coherent text representations. In contrast, TurnGuide aggregates these text tokens, creating a more cohesive grouping that simplifies the learning process for text generation, which leads to notable performance improvements. Second, Moshi TS fails to mitigate the inherently fragmented nature of speech semantics. Here, we propose a new way of understanding the semantic information within a speech audio. Speech audio combines multiple types of information together, including semantic, paralinguistic, and speaker information Cui et al. (2024), and the semantics within speech can be defined by its transcription, with speech containing semantic information only at the temporal points when each word is sounded. This fragmented nature increases the difficulty for coherent speech generation. Although Moshi TS incorporates text into the speech generation process, it inserts text tokens strictly at the time they are sounded in the speech. This approach only supplements speech with semantic information at points where semantic information already exists, offering limited potential for enhancing semantic representation. By comparison, TurnGuide groups texts within the same turn and processes them collectively. This method introduces denser semantic information, effectively guiding the speech generation process to improve coherence. Third, Moshi TS heavily depends on precise word-level timestamps from the ASR systems. In practice, these timestamps are not always exact, often being slightly ahead or behind the actual sounding moment, which can disrupt text-speech alignment and degrade performance. TurnGuide, however, leverages only the start time of each text chunk, significantly reducing reliance on word-level timestamp accuracy. This simpler alignment strategy not only lowers the risk of errors but also contributes to stronger overall performance.
Appendix D Corpus-Level Turn-Taking Evaluations
We also incorporate the results of corpus-level turn-taking evaluations. Following prior works, the turn-taking events considered here include Inter-Pausal Units (IPUs), pause, gap, and overlap (Nguyen et al., 2023). Figure 6 provides examples of these turn-taking events. To give a brief explanation of how these turn-taking events are extracted from the generated dialogue: First, the speech segments within each speaker channel are identified using a pyannote VAD module. An IPU is defined as continuous speech within a single speaker channel. Multiple speech segments are grouped into a single IPU if the pause between them is shorter than $\tau_{\text{IPU}}$. As previously mentioned, $\tau_{\text{ IPU }}$ is initially set to 0.5 seconds in the multi-modal turn segmentation and alignment framework (Section 3.1). However, for the evaluations, it is adjusted to 0.2 seconds in accordance with prior studies Nguyen et al. (2023); Veluri et al. (2024); Wang et al. (2025). A Pause refers to the absence of speech within the same speaker channel, while a Gap indicates the discontinuation of speech across different channels. Overlap refers to the time period during which both speakers are speaking simultaneously.
We calculate the Pearson correlation of the occurrence, cumulative duration, and average duration of the turn-taking events between the 30-second prompts and the corresponding 90-second continuations, as described in Veluri et al. (2024); Nguyen et al. (2023). This measures whether the model can generate dialogues with similar conversational features to the prompt. The metric is averaged across the four turn-taking events.
The corpus-level turn-taking evaluation results on the Fisher testset are presented in Table 7. We conduct evaluations under both unconditional and conditional settings, comparing TurnGuide with various baseline approaches. The results indicate that TurnGuide demonstrates comparable performance to the baselines in terms of Pearson correlation for occurrence, cumulative duration, and average duration. This finding suggests that TurnGuide, which integrates additional text streams into the modeling process, does not exhibit significant behavioral differences at the corpus level.
Remark 2
We include the corpus-level turn-taking results for the sake of completeness. However, as noted in Section 4.2.2, corpus-level statistics are less interpretable and not as suitable for evaluating the turn-taking abilities of FD-SLMs. Therefore, we recommend readers focus on the more reliable results assessed using the Full-Duplex-Bench benchmark (Section 4.2.2).
Appendix E The Fisher Dataset Split
The Fisher dataset is frequently referenced and employed in e2e FD-SLM studies. However, to date, there is no official training, validation, and test set split for the dataset, nor do existing papers provide details of their dataset splits. This poses a significant challenge for future research, as it becomes difficult to replicate the reported model performance in original studies. Specifically, when researchers use their own dataset splits to train models and subsequently compare their model performance with publicly available checkpoints from prior work, there is a risk of test set leakage. This occurs when an existing model has been trained on test samples included within a new split defined by researchers. To tackle this issue, we propose releasing our train/validation/test splits for the Fisher dataset and strongly encourage future studies to adopt these splits for a more equitable and reliable basis of comparison. Specifically, Fisher separates all the audio into 117 parts (000-116). We split the train/validation/test set using part IDs 000-110/111-113/114-116, and we randomly select roughly 1,000 samples in our Fisher test set to perform all the evaluations reported in this paper.
Appendix F Latency Analysis
In this section, we theoretically analyze the overall first-package latency of TurnGuide, which represents the time taken for the user to hear the first output audio sample. Since our model is a fine-tuned version of GLM-4-Voice, much of the latency calculation can refer to the original GLM-4-Voice paper. Specifically, the e2e latency is divided into three components: Speech Tokenization, LLM processing (including pre-filling and token decoding), and Speech Decoding.
With TurnGuide processing speech in chunks, the first-package latency corresponds to the total processing time (Speech Tokenization + LLM + Speech Decoding) for the first chunk. Using a speech chunk size of five tokens, the latency is calculated as follows: 1) The Speech Tokenization time for 5 tokens ($T_{st}$) is 0.01s, the LLM processing time ($T_{LLM}$), which includes pre-filling 5 user speech tokens along with special tokens and decoding a total of 10 assistant tokens (5 text tokens + 5 speech tokens), is 0.3s, and the Speech Decoding time ($T_{SD}$) for 10 speech tokens is 0.75s. Thus, the total latency is $T_{ALL}=T_{st}+T_{LLM}+T_{SD}=1.05s$. It’s worth noting that the optimal speech decoding process involves decoding only 5 speech tokens instead of 10. However, the original GLM-4-Voice vocoder requires at least 10 tokens to decode, meaning the vocoder’s latency can be further reduced in an optimized solution.
Appendix G Semantic Evaluation Prompts
We give prompts for the three settings of the GPT-score semantic evaluation used in Table 2: We use the prompt in Figure 7 for the unconditional setting, and use the prompt in Figure 8 for the two conditional settings.
Appendix H Explanations of Turn-Taking Events in Full-Duplex-Bench
This section details the four turn-taking events evaluated in the Full-Duplex-Bench. Specifically, 1) Smooth Turn-taking evaluates the model’s ability to seamlessly take over the conversation after the user’s speaking turn. 2) User Interruption assesses the model’s capability to recognize and adapt to new content introduced by the user during the interaction. 3) Pause Handling determines whether the model can remain silent during brief pauses by the user before they complete their speaking turn. 4) Backchanneling measures the model’s ability to provide brief acknowledgments at appropriate moments while the user is speaking. For a detailed explanations of the evaluation metrics, please refer to their original paper (Lin et al., 2025).
Appendix I Potential Risks
The primary ethical concern regarding TurnGuide involves safety and alignment risks arising from the model’s training data and the absence of post-hoc behavioral tuning. Specifically, since the model is trained on the Fisher dataset—a collection of real-world telephone conversations—it may inherit and replicate biased, sensitive, or harmful topics present in authentic human dialogue. These risks are compounded by the fact that the current iteration has not undergone Reinforcement Learning from Human Feedback (RLHF) (Lambert, 2025; Ouyang et al., 2022; Rafailov et al., 2023) or similar safety alignment procedures to ensure its outputs remain consistent with human preferences. Without a robust refusal mechanism or a safety-oriented backbone, the model’s enhanced capabilities for real-time interruption and overlapping speech could potentially be exploited to generate deceptive or inappropriate content. Consequently, TurnGuide is intended strictly for research purposes and is not suitable for production deployment without the integration of rigorous safety-tuning frameworks and content moderation safeguards.
Please evaluate the following two-speaker dialogue transcript for how meaningful the speech is (based on its content), focusing primarily on the part from 30 seconds to 120 seconds of the transcript. Use the following scale:
0: Completely meaningless; no coherent sentences, random words, or unintelligible.
0.5: Almost no meaning; isolated words or phrases, but no understandable ideas.
1: Extremely low meaning; rare, vague fragments of ideas, but mostly incoherent or off-topic.
1.5: Very little meaning; a few short, unclear ideas, but mostly disjointed or confusing.
2: Low meaning; some recognizable ideas or topics, but mostly unclear, incomplete, or off-topic.
2.5: Somewhat low meaning; a few coherent points, but overall lacks clarity or logical flow.
3: Moderate meaning; general topic is understandable, but there are gaps, unclear parts, or weak connections.
3.5: Fairly meaningful; mostly coherent and relevant, but with some confusion, repetition, or lack of detail.
4: Meaningful; clear and logical, with relevant and connected ideas, though may lack depth or detail.
4.5: Very meaningful; almost fully coherent, with well-developed, relevant, and connected ideas.
5: Extremely meaningful; highly coherent, clear, and detailed, with all ideas well connected and relevant.
Only output the final score (0, 0.5, 1, 1.5, ..., 5) **ONLY** according to the above rubric. Do not output anything else.
Figure 7: The prompt for GPT-score semantic evaluation under the unconditional setting.
Please evaluate the following two-speaker dialogue transcript for how meaningful the speech is (based on its content), only focusing on the model channels output and from 30 seconds to 120 seconds of the transcript. Use the following scale:
0: Completely meaningless; no coherent sentences, random words, or unintelligible.
0.5: Almost no meaning; isolated words or phrases, but no understandable ideas.
1: Extremely low meaning; rare, vague fragments of ideas, but mostly incoherent or off-topic.
1.5: Very little meaning; a few short, unclear ideas, but mostly disjointed or confusing.
2: Low meaning; some recognizable ideas or topics, but mostly unclear, incomplete, or off-topic.
2.5: Somewhat low meaning; a few coherent points, but overall lacks clarity or logical flow.
3: Moderate meaning; general topic is understandable, but there are gaps, unclear parts, or weak connections.
3.5: Fairly meaningful; mostly coherent and relevant, but with some confusion, repetition, or lack of detail.
4: Meaningful; clear and logical, with relevant and connected ideas, though may lack depth or detail.
4.5: Very meaningful; almost fully coherent, with well-developed, relevant, and connected ideas.
5: Extremely meaningful; highly coherent, clear, and detailed, with all ideas well connected and relevant.
Only output the final score (0, 0.5, 1, 1.5, ..., 5) **ONLY** according to the above rubric. Do not output anything else.
Figure 8: The prompt for GPT-score semantic evaluation under the two conditional settings.
## Pairwise comparison instructions
### Task overview
You will be shown **30 pairs** of two-speaker dialogue samples generated by different systems.
For each pair, your job is to choose **which dialogue has better semantic quality** (i.e., which one is more meaningful and coherent in what the speakers say).
### What you will receive for each pair
For each pair, you will be given:
- **Audio** of Dialogue A and Dialogue B.
- **Transcripts** for both dialogues (automatically generated with Whisper-large-v3 using the stable-ts package).
The order (A/B) is **randomized** for every pair; do not assume A or B is better.
### What "better semantic quality" means
When deciding which dialogue is better, focus on **content meaning and coherence**, not just fluency.
Prefer the dialogue that is:
- More **understandable** and stays on a recognizable topic.
- More **coherent**, with logical connections between utterances.
- Less **contradictory**, less nonsensical, and less off-topic.
- More **informative/complete** rather than fragmented or empty.
### What to ignore (or de-emphasize)
Do **not** primarily judge based on:
- Audio naturalness (voice quality, prosody, accent, loudness) unless it prevents understanding.
- Turn-taking style (interruptions/overlap) unless it makes the content hard to follow.
- Minor transcription errors, as long as the intended meaning is still clear from audio + context.
### How to make your choice
For each pair, listen to both dialogues (you may replay) and read transcripts as needed.
Then select exactly one option:
- **A is better** (Dialogue A has higher semantic quality).
- **B is better** (Dialogue B has higher semantic quality).
### Consistency and independence
Judge each pair **independently**; do not try to keep scores balanced across the full study.
Make the best decision based only on the evidence in that pair (audio + transcript).
Figure 9: The instructions given to the participants in the pairwise comparison human evaluation.